# Energy exchange between electrons and ions in ion temperature gradient turbulence

### v1

@@ Abstract
Microturbulence in magnetic confined plasmas contributes to energy exchange between particles of different species as well as the particle and heat fluxes. Although the effect of turbulent energy exchange has not been considered significant in previous studies, it is anticipated to have a greater impact than collisional energy exchange in low collisional plasmas such as those in future fusion reactors. In this study, gyrokinetic simulations are performed to evaluate the energy exchange in ion temperature gradient (ITG) turbulence. The energy exchange due to the ITG turbulence mainly consists of the cooling of ions in the $\nabla B$-curvature drift motion and the heating of electrons streaming along a field line. It is found that the ITG turbulence transfers energy from ions to electrons regardless of whether the ions or electrons are hotter, which is in marked contrast to the energy transfer by Coulomb collisions. This implies that the ITG turbulence should be suppressed from the viewpoint of sustaining the high ion temperature required for fusion reactions since it prevents energy transfer from alpha-heated electrons to ions as well as enhancing ion heat transport toward the outside of the reactor. Furthermore, linear and nonlinear simulation analyses confirm the feasibility of quasilinear modeling for predicting the turbulent energy exchange in addition to the particle and heat fluxes.
@@ I Introduction
Numerous studies on anomalous transport of particles and heat generated by microscopic turbulence have been done, based on gyrokinetic theory and simulationAntonsen ; CTB ; F-C ; Dimits ; Krommes2012 ; Garbet2010 ; Idomura2006 ; Horton . However, there have been fewer theoretical works on energy exchange between different particle species due to turbulence,Sugama1996 ; Sugama2009 ; Waltz and a small number of simulations have been performed to investigate the effect of turbulent energy exchange on the evolution of temperature profiles. Candy In Ref. Candy , it is shown from gyrokinetic simulations that the effect of turbulent energy exchange is negligibly small under conditions of DIII-D shot 128913. However, sufficient comparative studies have not been made between collisional and turbulent energy exchanges in a wide range of conditions. In particular, in the case of high temperature plasmas, the impact of collisional energy exchange is expected to be small due to the low collision frequency, while turbulent energy exchange can work actively even in collisionless plasmas. Since collisional heat transfer from alpha-heated electrons to ions is expected to play a critical role in sustaining burning plasmas in future reactors, it is an important issue to compare the effects of turbulence and collisions on the energy exchange between ions and electrons in high temperature plasmas. In the present paper, we evaluate the energy exchange in ion temperature gradient (ITG) turbulence by gyrokinetic turbulence simulations and investigate its properties, such as dependence on the ratio between ion and electron temperatures in comparison to those of the collisional energy exchange.
@@
To perform simulations that predict global density and temperature profiles, it is practical to use turbulent transport models such as quasilinear ones,Casati ; Bourdelle ; Staebler ; Narita ; Toda ; Parker rather than running direct turbulence simulations for all cases. In fact, it has been shown that these models can reproduce particle and heat fluxes obtained from gyrokinetic turbulence simulations within acceptable errors. A quasilinear model for turbulent energy exchange is shown in Ref. Barnes where an electron drift wave instability is treated in a slab geometry with no temperature gradient. Detailed studies have not been conducted on modeling turbulent energy exchange under more complex conditions such as those for toroidal ITG mode remains. Quasilinear models are based on the assumption that the ratios of turbulent transport fluxes and energy exchange to the squared potential fluctuation amplitude estimated by linear analyses, which are called quasilinear weightsStaebler , take approximately the same values as those ratios in a steady state of turbulence obtained by nonlinear simulations. In this work, linear and nonlinear simulation results are compared to show the validity of this assumption in the quasilinear modeling of energy exchange as well as particle and heat transport fluxes in the ITG turbulence. We also discuss physical mechanisms and the quasilinear modeling of turbulent energy exchange by wavenumber spectral analyses of entropy balance in microturbulenceSugama1996 ; Sugama2009 ; Nakata ; Maeyama .
@@ II.1 Gyrokinetic equations
The plasma distribution function $F_{a}$ can be written as $F_{a}(\bm{x},\bm{v},t)=f_{a}(\bm{x},\bm{v},t)+\widehat{f}_{a}(\bm{x},\bm{v},t)$, where $f_{a}$ and $\widehat{f}_{a}$ represent the ensemble-averaged and fluctuation parts, respectively, and subscript $a$ denotes the particle species. The space-time scales of variations in the ensemble-averaged part $f_{a}$ are much larger than those of the fluctuation part $\widehat{f}_{a}$ so that the ensemble average can also be regarded as the local space–time average. In the gyrokinetic theory, perturbations such as $\widehat{f}_{a}$ are assumed to satisfy the gyrokinetic ordering,
@@
where $\widehat{\phi}$ and $\widehat{\bm{B}}$ are the perturbed electrostatic potential and the perturbed magnetic field, respectively. Here, $\Omega_{a}=e_{a}B/m_{a}c$ , $\rho_{ta}=v_{ta}/\Omega_{a}$ , $m_{a}$, $e_{a}$, $T_{a}$, and $v_{ta}\equiv\sqrt{T_{a}/m_{a}}$ are the gyrofrequency, thermal gyroradius, mass, charge, temperature, and thermal velocity, respectively. The characteristic wavenumbers (in directions parallel and perpendicular to the background magnetic field $\bm{B}$), frequency, and equilibrium scale length are represented by $k_{\parallel}$, $k_{\perp}$, $\omega$, and $L$, respectively.
@@
where $\bm{\rho}_{a}=\bm{b}\times\bm{v}/\Omega_{a}$ with $\bm{b}=\bm{B}/|\bm{B}|$. Here, $h_{a\bm{k}_{\perp}}$ represents the nonadiabatic distribution function which is calculated by the gyrokinetic equation,
$\left(\frac{\partial}{\partial t}+v_{\parallel}\bm{b}\cdot\nabla+i\bm{k}_{\perp}\cdot(\bm{v}_{E}+\bm{v}_{da})\right)h_{a\bm{k}_{\perp}}$ (5)
@@
$+\frac{c}{B}\sum_{\bm{k}_{\perp}^{\prime}+\bm{k}_{\perp}^{\prime\prime}=\bm{k}_{\perp}}\left[\bm{b}\cdot\left(\bm{k}_{\perp}^{\prime}\times\bm{k}_{\perp}^{\prime\prime}\right)\right]\psi_{a\bm{k}_{\perp}^{\prime}}h_{a\bm{k}_{\perp}^{\prime\prime}}+C_{a},$
where $v_{\parallel}=\bm{v}\cdot\bm{b}$, $\bm{v}_{E}=c\bm{E}\times\bm{b}/B$, $\bm{v}_{da}=c\bm{b}\times(\mu\nabla B+m_{a}v_{\parallel}^{2}\bm{b}\cdot\nabla\bm{b})/e_{a}B$, $\bm{v}_{*a}=cT_{a}\bm{b}\times\left\{\nabla\ln n_{a}+\left(w-3/2\right)\nabla\ln T_{a}\right\}/e_{a}B$, $C_{a}$ is the gyrokinetic collision term, and $\bm{E}=-\nabla\Phi$ is the background electric field. In Eq. (5), $h_{a\bm{k}_{\perp}}$ is regarded as a function of time $t$ and phase space variables ($\bm{x}$, $w=m_{a}v^{2}/2$, $\mu=m_{a}v_{\perp}^{2}/2B$). The gyrophase-averaged perturbed potential function $\psi_{a\bm{k}_{\perp}}$ is defined in terms of the perturbed electrostatic potential $\phi_{\bm{k}_{\perp}}$ and the perturbed vector potential $\bm{A}_{\bm{k}_{\perp}}$ as
$\psi_{a\bm{k}_{\perp}}$ $=$ $\oint\frac{d\xi}{2\pi}\exp\left(i\bm{k}_{\perp}\cdot\bm{\rho}_{a}\right)\left(\phi_{\bm{k}_{\perp}}-\frac{\bm{v}}{c}\cdot\bm{A}_{\bm{k}_{\perp}}\right)$ (6)
@@
where $\lambda_{D}=1/\sqrt{\sum_{a}{4\pi n_{a}e_{a}^{2}/T_{a}}}$ is the Debye length. The time evolution of nonadiabatic distribution function $h_{a\bm{k}_{\perp}}$ can be obtained by solving Eq.(5) combined with Eqs. (7)–(9).
@@ II.2 Energy and entropy balance equations
We now consider two balance equations related to the perturbed entropy density and the energy. These valuables are associated with the particle and heat fluxes and the energy exchange between electrons and ions. The entropy density due to turbulence $\delta S_{a}$ is defined by the difference between the macroscopic entropy density $S_{Ma}\equiv-\int d^{3}vf_{a}\log f_{a}$ and the ensemble averaged microscopic entropy density $\langle S_{ma}\rangle_{\rm ens}\equiv-\left\langle\int d^{3}v\left(f_{a}+\widehat{f}_{a}\right)\log\left(f_{a}+\widehat{f}_{a}\right)\right\rangle_{\rm ens}$ as
@@
where *** denotes the complex conjugate and $n_{a\bm{k}_{\perp}}\equiv\int{d^{3}vf_{a\bm{k}_{\perp}}}$ is the density perturbation with the perpendicular wavenumber vector $\bm{k}_{\perp}$. Multiplying Eq. (5) by $h_{a\bm{k}_{\perp}}^{*}/f_{Ma}$ and taking the ensemble and flux-surface averages, we can derive the entropy balance equation,
$\frac{\partial}{\partial t}\left(\delta S_{ha}\right)$ $=$ $\frac{\Gamma^{\rm turb}_{a}}{L_{pa}}+\frac{q^{\rm turb}_{a}}{T_{a}L_{Ta}}+\frac{Q^{\rm turb}_{a}}{T_{a}}+D_{a}$
@@
$\sum_{a}{Q^{\rm turb}_{a}}=\sum_{a}\sum_{\bm{k}_{\perp}}\frac{e_{a}}{2}\frac{\partial}{\partial t}\operatorname{\rm{Re}}\Bigg{\langle}\Bigg{\langle}\int d^{3}vh_{a\bm{k}_{\perp}}^{*}\psi_{a\bm{k}_{\perp}}\Bigg{\rangle}\Bigg{\rangle}$
$=\frac{1}{8\pi}\frac{\partial}{\partial t}\sum_{\bm{k}_{\perp}}\Bigg{\langle}\Bigg{\langle}\left(k_{\perp}^{2}+\lambda_{D}^{-2}\right)|\phi_{\bm{k}_{\perp}}|^{2}-|\bm{B}_{\bm{k}_{\perp}}|^{2}\Bigg{\rangle}\Bigg{\rangle}.$ (16)
Using the transport time scale ordering $\partial\langle\langle\cdots\rangle\rangle/\partial t=\mathcal{O}(\delta^{2})$, we obtain $\sum_{a}{Q^{\rm turb}_{a}}=0$ to the lowest order in $\delta$. Therefore, we can interpret $Q^{\rm turb}_{a}$ as the energy exchange between different plasma species. The last term in Eq. (II.2) $D_{a}$ represents collisional dissipation written as
$D_{a}=\sum_{\bm{k}_{\perp}}\operatorname{\rm{Re}}\Bigg{\langle}\Bigg{\langle}\int d^{3}v\frac{f_{a\bm{k}_{\perp}}^{*}}{f_{Ma}}C_{a}\Bigg{\rangle}\Bigg{\rangle}$ (17)
The left-hand side of Eq. (II.2) vanishes in the steady state of turbulence where the entropy production due to turbulent transport and heating balance with the collisional dissipation.
@@
$\frac{\partial}{\partial t}\left(\frac{3}{2}p_{a}\right)+\frac{1}{V^{\prime}}\frac{\partial}{\partial r}\left[V^{\prime}\left(q_{a}+\frac{5}{2}T_{s}\Gamma_{a}\right)\right]$ (18)
$=\frac{(\Gamma_{a}^{\rm{cl}}+\Gamma_{a}^{\rm{ncl}})}{n_{a}}\frac{\partial p_{a}}{\partial r}+\langle\bm{u}_{a}\cdot(\nabla\cdot\bm{\pi}_{a})\rangle_{a}+Q^{\rm coll}_{a}$
$\hskip 14.22636pt-e_{a}\Gamma_{a}^{\rm{turb}}\frac{\partial\Phi}{\partial r}+Q^{\rm turb}_{a}$
where $\langle\cdots\rangle_{a}$ represents the flux-surface average, and $\bm{\pi}_{s}$ is the viscosity tensor. The radial particle and heat fluxes denoted by $\Gamma_{a}$ and $q_{a}$, respectively, are written as
@@
where $\nu_{e}$ is the electron-ion collision frequency, ${\bf R}_{e}$ the collisional friction force for electrons, and $({\bf u}_{e}-{\bf u}_{i})$ the difference between the electron and ion flow velocities. We see that the collisional energy transfer $Q^{\rm coll}_{i}$ from electrons to ions is proportional to the product of the collision frequency and the temperature difference between electrons and ions.
The sum of the last two terms on the right-hand side of Eq.(18) can be rewritten asSugama1996
$-e_{a}\Gamma_{a}^{\rm{turb}}\frac{\partial\Phi}{\partial r}+Q^{\rm turb}_{a}$
[…]
As pointed out in Ref.Sugama1996 , the background radial electric field $E_{r}=-\partial\Phi/\partial r$ enters the gyrokinetic equation, Eq. (5) only in the form of the Doppler shift $(\partial/\partial t+i\bm{k}_{\perp}\cdot\bm{v}_{E})$ and does not appear explicitly in Eqs. (6)–(9). Therefore, for the solutions of Eq. (5) and Eqs. (7)–(9), $e^{i\bm{k}_{\perp}\cdot\bm{v}_{E}}h_{a\bm{k}_{\perp}}$ and $e^{i\bm{k}_{\perp}\cdot\bm{v}_{E}}\psi_{a\bm{k}_{\perp}}$ are independent of $E_{r}$. Then, even though each of the first and second terms on the left-hand side of Eq. (II.2) depends on $E_{r}$, their sum does not. In the present paper, we evaluate $Q^{\rm turb}_{a}$ by the gyrokinetic simulation for $E_{r}=0$ although we should note that the results of $Q^{\rm turb}_{a}$ for $E_{r}=0$ are equivalent to those of $e_{a}\Gamma_{a}^{\rm{turb}}E_{r}+Q^{\rm turb}_{a}$ for any $E_{r}$. We also find from Eq. (14) that the right-hand side of the entropy balance equation, Eq. (II.2), contains this sum of the terms $e_{a}\Gamma_{a}^{\rm{turb}}E_{r}+Q^{\rm turb}_{a}$ which takes a value independent of $E_{r}$.

### v2

@@ Abstract
Microturbulence in magnetic confined plasmas contributes to energy exchange between particles of different species as well as the particle and heat fluxes. Although the effect of turbulent energy exchange has not been considered significant in previous studies, it is anticipated to have a greater impact than collisional energy exchange in low collisional plasmas such as those in future fusion reactors. In this study, gyrokinetic simulations are performed to evaluate the energy exchange due to ion temperature gradient (ITG) turbulence in a tokamak configuration. The energy exchange due to the ITG turbulence mainly consists of the cooling of ions in the $\nabla B$-curvature drift motion and the heating of electrons streaming along a field line. It is found that the ITG turbulence transfers energy from ions to electrons regardless of whether the ions or electrons are hotter, which is in marked contrast to the energy transfer by Coulomb collisions. This implies that the ITG turbulence should be suppressed from the viewpoint of sustaining the high ion temperature required for fusion reactions since it prevents energy transfer from alpha-heated electrons to ions as well as enhancing ion heat transport toward the outside of the reactor. Furthermore, linear and nonlinear simulation analyses confirm the feasibility of quasilinear modeling for predicting the turbulent energy exchange in addition to the particle and heat fluxes.
@@ I Introduction
Numerous studies on anomalous transport of particles and heat generated by microscopic turbulence have been done, based on gyrokinetic theory and simulationAntonsen ; CTB ; F-C ; Dimits ; Krommes2012 ; Garbet2010 ; Idomura2006 ; Horton . However, there have been fewer theoretical works on energy exchange between different particle species due to turbulence,Sugama1996 ; Sugama2009 ; Hinton ; Waltz and a small number of simulations have been performed to investigate the effect of turbulent energy exchange on the evolution of temperature profiles. Candy In Ref. Candy , it is shown from gyrokinetic simulations that the effect of turbulent energy exchange is negligibly small under conditions of DIII-D shot 128913. However, sufficient comparative studies have not been made between collisional and turbulent energy exchanges in a wide range of conditions. In particular, in the case of high temperature plasmas, the impact of collisional energy exchange is expected to be small due to the low collision frequency, while turbulent energy exchange can work actively even in collisionless plasmas. As examples of other works related to turbulent energy exchange, there have been studies on the scaling of the turbulent transport and heating of impurities in magnetized plasmasBarnes_scaling and the thermal equilibration of ions and electrons by turbulence in astrophysical plasmasKawazura .
Since collisional heat transfer from alpha-heated electrons to ions is expected to play a critical role in sustaining burning plasmas in future reactors, it is an important issue to compare the effects of turbulence and collisions on the energy exchange between ions and electrons in high temperature plasmas. In the present paper, we evaluate the energy exchange in ion temperature gradient (ITG) turbulence by gyrokinetic turbulence simulations and investigate its properties, such as dependence on the ratio between ion and electron temperatures in comparison to those of the collisional energy exchange.
@@
To perform simulations that predict global density and temperature profiles, it is practical to use turbulent transport models such as quasilinear ones,Casati ; Bourdelle ; Staebler ; Narita ; Toda ; Parker rather than running direct turbulence simulations for all cases. In fact, it has been shown that these models can reproduce particle and heat fluxes obtained from gyrokinetic turbulence simulations within acceptable errors. A quasilinear model for turbulent energy exchange is shown in Ref. Barnes where an electron drift wave instability is treated in a slab geometry with no temperature gradient. Detailed studies have not been conducted on modeling turbulent energy exchange under more complex conditions such as those for toroidal ITG mode remains. Quasilinear models are based on the assumption that the ratios of turbulent transport fluxes and energy exchange to the squared potential fluctuation amplitude estimated by linear analyses, which are called quasilinear weightsStaebler , take approximately the same values as those ratios in a steady state of turbulence obtained by nonlinear simulations. In this work, linear and nonlinear simulation results are compared to show the validity of this assumption in the quasilinear modeling of energy exchange as well as particle and heat transport fluxes in the ITG turbulence. We here note that this work demonstrates only the feasibility of quasilinear modeling but does not present a saturation rule which is necessary for developing a quasilinear model. We also discuss physical mechanisms and the quasilinear modeling of turbulent energy exchange by wavenumber spectral analyses of entropy balance in microturbulenceSugama1996 ; Sugama2009 ; Nakata ; Maeyama .
@@ II.1 Gyrokinetic equations
The plasma distribution function $F_{a}$ of the position vector $\bm{x}$, velocity vector $\bm{v}$, and time $t$ can be written as $F_{a}(\bm{x},\bm{v},t)=f_{a}(\bm{x},\bm{v},t)+\widehat{f}_{a}(\bm{x},\bm{v},t)$, where $f_{a}$ and $\widehat{f}_{a}$ represent the ensemble-averaged and fluctuation parts, respectively, and subscript $a$ denotes the particle species. The space-time scales of variations in the ensemble-averaged part $f_{a}$ are much larger than those of the fluctuation part $\widehat{f}_{a}$ so that the ensemble average can also be regarded as the local space–time average. In the gyrokinetic theory, perturbations such as $\widehat{f}_{a}$ are assumed to satisfy the gyrokinetic ordering,
@@
where $\widehat{\phi}$ and $\widehat{\bm{B}}$ are the perturbed electrostatic potential and the perturbed magnetic field, respectively. Here, $\Omega_{a}=e_{a}B/m_{a}c$ , $\rho_{ta}=v_{ta}/\Omega_{a}$ , $m_{a}$, $e_{a}$, $T_{a}$, $\bm{B}$, $c$, and $v_{ta}\equiv\sqrt{T_{a}/m_{a}}$ are the gyrofrequency, thermal gyroradius, mass, charge, temperature, background magnetic field, speed of light, and thermal velocity, respectively. The characteristic wavenumbers (in directions parallel and perpendicular to the background magnetic field), frequency, and equilibrium scale length are represented by $k_{\parallel}$, $k_{\perp}$, $\omega$, and $L$, respectively.
@@
where $\bm{\rho}_{a}=\bm{b}\times\bm{v}/\Omega_{a}$ with $\bm{b}=\bm{B}/B$ and $B=|\bm{B}|$. Here, $h_{a\bm{k}_{\perp}}$ represents the nonadiabatic distribution function which is calculated by the gyrokinetic equation,
$\left(\frac{\partial}{\partial t}+v_{\parallel}\bm{b}\cdot\nabla+i\bm{k}_{\perp}\cdot(\bm{v}_{E}+\bm{v}_{da})\right)h_{a\bm{k}_{\perp}}$
@@
$+\frac{c}{B}\sum_{\bm{k}_{\perp}^{\prime}+\bm{k}_{\perp}^{\prime\prime}=\bm{k}_{\perp}}\left[\bm{b}\cdot\left(\bm{k}_{\perp}^{\prime}\times\bm{k}_{\perp}^{\prime\prime}\right)\right]\psi_{a\bm{k}_{\perp}^{\prime}}h_{a\bm{k}_{\perp}^{\prime\prime}}+C_{a}^{GK},$ (5)
where $v_{\parallel}=\bm{v}\cdot\bm{b}$, $\bm{v}_{E}=c\bm{E}\times\bm{b}/B$, $\bm{v}_{da}=c\bm{b}\times(\mu\nabla B+m_{a}v_{\parallel}^{2}\bm{b}\cdot\nabla\bm{b})/e_{a}B$, $\bm{v}_{*a}=cT_{a}\bm{b}\times\left\{\nabla\ln n_{a}+\left(w-3/2\right)\nabla\ln T_{a}\right\}/e_{a}B$, $C_{a}^{GK}$ is the gyrokinetic collision termSugama2009 , and $\bm{E}=-\nabla\Phi$ is the background electric field. Here, the background $E\times B$ flow is assumed to be $\bm{v}_{E}\sim\delta v_{ti}$, and the gradient scale length of $\bm{v}_{E}$ is estimated as $L$. Therefore, the effect of electric field shear is neglected. The detail is discussed in Appendix A. In Eq. (II.1), $h_{a\bm{k}_{\perp}}$ is regarded as a function of time $t$ and phase space variables ($\bm{x}$, $w=m_{a}v^{2}/2$, $\mu=m_{a}v_{\perp}^{2}/2B$), where $v=|\bm{v}|$ and $v_{\perp}=|\bm{v}-v_{\parallel}\bm{b}|$. The gyrophase-averaged perturbed potential function $\psi_{a\bm{k}_{\perp}}$ is defined in terms of the perturbed electrostatic potential $\phi_{\bm{k}_{\perp}}$ and the perturbed vector potential $\bm{A}_{\bm{k}_{\perp}}$ as
$\psi_{a\bm{k}_{\perp}}$ $=$ $\oint\frac{d\xi}{2\pi}\exp\left(i\bm{k}_{\perp}\cdot\bm{\rho}_{a}\right)\left(\phi_{\bm{k}_{\perp}}-\frac{\bm{v}}{c}\cdot\bm{A}_{\bm{k}_{\perp}}\right)$
@@
where $\lambda_{D}=1/\sqrt{\sum_{a}{4\pi n_{a}e_{a}^{2}/T_{a}}}$ is the Debye length. The time evolution of nonadiabatic distribution function $h_{a\bm{k}_{\perp}}$ can be obtained by solving Eq.(II.1) combined with Eqs. (7)–(9).
@@ II.2 Energy and entropy balance equations
We now consider two balance equations related to the perturbed entropy density and the energy. These variables are associated with the particle and heat fluxes and the energy exchange between electrons and ions. The entropy density due to turbulence $\delta S_{a}$ is defined by the difference between the macroscopic entropy density $S_{Ma}\equiv-\int d^{3}vf_{a}\log f_{a}$ and the ensemble averaged microscopic entropy density $\langle S_{ma}\rangle_{\rm ens}\equiv-\left\langle\int d^{3}v\left(f_{a}+\widehat{f}_{a}\right)\log\left(f_{a}+\widehat{f}_{a}\right)\right\rangle_{\rm ens}$ as
@@
where * denotes the complex conjugate and $n_{a\bm{k}_{\perp}}\equiv\int{d^{3}vf_{a\bm{k}_{\perp}}}$ is the density perturbation with the perpendicular wavenumber vector $\bm{k}_{\perp}$. Multiplying Eq. (II.1) by $h_{a\bm{k}_{\perp}}^{*}/f_{Ma}$ and taking the ensemble and flux-surface averages, we can derive the entropy balance equation,
$\frac{\partial}{\partial t}\left(\delta S_{ha}\right)$ $=$ $\frac{\Gamma^{\rm turb}_{a}}{L_{pa}}+\frac{q^{\rm turb}_{a}}{T_{a}L_{Ta}}+\frac{Q^{\rm turb}_{a}}{T_{a}}+D_{a}$ (12)
@@
$\sum_{a}{Q^{\rm turb}_{a}}=\frac{1}{8\pi}\frac{\partial}{\partial t}\sum_{\bm{k}_{\perp}}\Bigg{\langle}\Bigg{\langle}\left(k_{\perp}^{2}+\lambda_{D}^{-2}\right)|\phi_{\bm{k}_{\perp}}|^{2}-|\bm{B}_{\bm{k}_{\perp}}|^{2}\Bigg{\rangle}\Bigg{\rangle},$
where $\bm{B}_{\bm{k}_{\perp}}=i\bm{k}_{\perp}\times\bm{A}_{\bm{k}_{\perp}}$. Using the transport time scale ordering $\left(\partial\langle\langle\cdots\rangle\rangle/\partial t\right)/\langle\langle\cdots\rangle\rangle=\mathcal{O}(\delta^{2})$, we obtain $\sum_{a}{Q^{\rm turb}_{a}}=0$ to the lowest order in $\delta$. Therefore, we can interpret $Q^{\rm turb}_{a}$ as the energy exchange between different plasma species. The last term in Eq. (12) $D_{a}$ represents collisional dissipation written as
$D_{a}\equiv\sum_{\bm{k}_{\perp}}D_{a\bm{k}_{\perp}}=\sum_{\bm{k}_{\perp}}\operatorname{\rm{Re}}\Bigg{\langle}\Bigg{\langle}\int d^{3}v\frac{h_{a\bm{k}_{\perp}}^{*}}{f_{Ma}}C_{a}^{GK}\Bigg{\rangle}\Bigg{\rangle}$ (17)
The left-hand side of Eq. (12) vanishes in the steady state of turbulence where the entropy production due to turbulent transport and heating balance with the collisional dissipation.
@@
$\frac{\partial}{\partial t}\left(\frac{3}{2}p_{a}\right)+\frac{1}{V^{\prime}}\frac{\partial}{\partial r}\left[V^{\prime}\left(q_{a}+\frac{5}{2}T_{a}\Gamma_{a}\right)\right]$
$=\frac{(\Gamma_{a}^{\rm{cl}}+\Gamma_{a}^{\rm{ncl}})}{n_{a}}\frac{\partial p_{a}}{\partial r}+\langle\bm{u}_{a}\cdot(\nabla\cdot\bm{\pi}_{a})\rangle+Q^{\rm coll}_{a}$
$\hskip 14.22636pt-e_{a}\Gamma_{a}^{\rm{turb}}\frac{\partial\Phi}{\partial r}+Q^{\rm turb}_{a}$ (18)
[…]
@@
where $\nu_{e}$ is the electron-ion collision frequency, $\ln\Lambda$ is the Coulomb logarithm, ${\bf R}_{e}$ the collisional friction force for electrons, and $({\bf u}_{e}-{\bf u}_{i})$ the difference between the electron and ion flow velocities. We see that the collisional energy transfer $Q^{\rm coll}_{i}$ from electrons to ions is proportional to the product of the collision frequency and the temperature difference between electrons and ions.
Appendix A shows that in the present model, the effect of the radial electric field $E_{r}$ on turbulent energy exchange appears only in the Doppler shift. The sum of the last two terms on the right-hand side of Eq. (II.2) is essentially equivalent to turbulent energy exchange Eq. (15) in $E_{r}=0$. Therefore, we henceforth discuss turbulent energy exchange with $E_{r}=0$.